QUESTION 13

On September 1, 2011, a company purchased a weaving machine for $239,800. The machine has an estimated useful life of 8 years and an estimated residual value of $17,800. Additionally, it is estimated that the machine would produce 740,000 bolts of woven fabric over its useful life. The company ended up selling the machine in 2018 after 1 month of use.

The following budgeted and actual activity levels were provided to support your work:


Assuming the company uses the straight line depreciation method, select the answer below that shows the machine's net book value at 12/31/2013. If applicable, round your intermediate calculations to 3 decimals and your final answer to the nearest $1.

A.

$156,550

B.

$138,750

C.

$160,000

D.

$175,050

E.

$157,250
Answer #1

Answer:-

D) 175050

DEPRECIATION = COST - SALVAGE VALUE / YEARS

= 239800 - 17800 / 8 = 27750

Given time 2 years 4 months

1september 2011 tp 12/31/2013

So depreciation = 27750 ×2 + (27750×4/12)

= 64750

Value = 239800 - 64750 = 175050